1. How many grams of O are present in 3.65 grams of dioxygen difluoride ?
------ grams O.
2. How many grams of dioxygen difluoride contain 4.01 grams of F ?
------- grams dioxygen difluoride.

3. How many grams of Co3+ are present in 1.13 grams of cobalt(III) bromide?
----- grams Co3+

4. How many grams of cobalt(III) bromide contain 2.36 grams of Co3+?
------- grams cobalt(III) bromide.
Answer #1

1. Mole = mass/molar mass

= 3.65/70=> 0.052 mol

Mass of oxygen = 2×0.052×16=> 1.67 g

2. Mole of F = 4.01/19 => 0.211 mol

Mass of dioxygen difluoride = 0.211×70/2

=> 7.39 g

3. Moles = 1.13/299 => 0.00378 mol

Mass of Co3+ = 0.000378×59 => 0.223 g

4. Mole = 2.36/59 => 0.04 mol

Mass of cobalt (lll) bromide = 0.04×299 => 11.96 g